a rubber ball of mass 50 g falls from a height of 1 m and rebounds to a height of 0.5 m. find the change in momentum of the ball

Grade:11

1 Answers

Arun
6 years ago

Initial velocity of ball at A = 0.
Final velocity of ball at B = v
                                    s = 1 m
As v2 - u2 = 2as
v2 - 02 = 2 x 9.8 x 1 or v = ?19.6 m/s
Let u' be the velocity of rebound of the ball.
s' = 50 cm = 0.5 m, g = - 9.8 m/s2
Final velocity at C = v' = 0
As v2 - u2 = 2as
Therefore 02 - (u')2 = 2 x (-9.8) x 0.5
or    u' = ?9.8 m/s
Now, Impulse =  Change in momentum
                      = mv - (-mu') = mv + mu' = m(v+ u)
                      = 50/1000 (?19.6 + ?9.8)
                      = 1/20 (4.427 + 3.130) = 0.378 Ns